The orbital angular momentum of a single 2s electron is (h is the
Planck’s constant) :
(A) h/2
(B) h/4
(C) h√ 2 / 2
(D) Zero
1
Expert's answer
2019-04-15T09:42:05-0400

Orbital angular momentum:


"L=\frac{\sqrt{l(l+1)}h}{2 \pi}"

For s-orbital:


"l=0"

Thus,


"L=0"

Answer: (D) Zero.